# MISS: A Generative Pretraining and Finetuning Approach for Med-VQA

### v1

@@ Abstract
Medical visual question answering (VQA) is a challenging multimodal task, where Vision-Language Pre-training (VLP) models can effectively improve the generalization performance. However, most methods in the medical field treat VQA as an answer classification task which is difficult to transfer to practical application scenarios. Additionally, due to the privacy of medical images and the expensive annotation process, large-scale medical image-text pairs datasets for pretraining are severely lacking. In this paper, we propose a large-scale MultI-task Self-Supervised learning based framework (MISS) for medical VQA tasks. Unlike existing methods, we treat medical VQA as a generative task. We unify the text encoder and multimodal encoder and align image-text features through multi-task learning. Furthermore, we propose a Transfer-and-Caption method that extends the feature space of single-modal image datasets using large language models (LLMs), enabling those traditional medical vision-field task data to be applied to VLP. Experiments show that our method achieves excellent results with fewer multimodal datasets and demonstrates the advantages of generative VQA models. The code and model weights will be released upon the paper’s acceptance.
Index Terms—  Medical visual question answering, vision-language pre-training, multi-modal learning
@@ 1 Introduction
Medical Visual Question Answering (Med-VQA) is a multi-modal task based on vision and language, aiming to provide corresponding answers to given images and questions. Compared with theVQA of natural images, Med-VQA requires a deeper and more accurate understanding of medical images. At the same time, due to the privacy of medical images and the high cost of high-quality text annotation, large-scale datasets for training Med-VQA are extremely scarce. Therefore, currently, Med-VQA is still a highly challenging task.
Thanks to the development of Convolutional Neural Networks (CNN) and Natural Language Processing (NLP) techniques, some works[16][15][17] have attempted to use CNN and Recurrent Neural Networks (RNN) to extract image and text features respectively for VQA tasks. With the emergence of transformers[7], image features and text features can be more easily embedded into the feature space with the same dimension, and VLP models[5] [6] have emerged continuously and have been proven to be effective solutions for downstream multi-modal tasks. While effective, these VLP models suffer from several limitations, such as domain gaps when applied to medical fields.
Currently, some multimodal models specialized in the medical domain have been proposed, such as M3AE [13], MRM [10], and CMITM [11], which unify masked autoencoders (MAE) [8] and masked language modeling (MLM) pre-training to learn joint representations of images and texts; MUMC [12] utilizes masked image modeling(MIM) by sending masked images to the image encoder as data augmentation; PMC-CLIP builds a new large medical dataset and trains it on a CLIP[6] model which pretrained on natural images. However, the above Med-VQA models still have two key problems:
@@
b. They utilize image-text pairs crawled from the article centres for pre-training, which contain a lot of noise, and high-quality open-source medical images for other tasks, such as classification and segmentation, have been ignored.
In this paper, we propose a new pre-training and fine-tuning paradigm for medical image-text tasks, named MISS, and apply it to the Med-VQA task. Unlike previous dual-tower multi-modal models, we treat Med-VQA as an answer-generating task, making our method directly applied to real-world scenarios and generating responses that more closely match human expression. We innovatively unify the text encoder and multi-modal encoder, building a Joint Text-Multimodal (JTM) encoder and enabling it to learn joint feature representations using a multi-task learning approach.
@@
$\bullet$ We introduce a new pretraining and fine-tuning framework named MISS. Not considering the Universal Large Vision-Language models (parameters more than 1B), it is the first pure generative VQA model in the medical field.
Fig. 1: Pretraining (a) and Finetuning (b) of our proposed method. We propose a pretraining and finetuning framework Miss for Med-VQA tasks which is composed of an image encoder, a JTM encoder, and a text decoder. ITC, ITM, and MLM Learning are used for pretraining. In the finetuning stage, the joint feature interacts with tokenized answers for MLM Learning.
@@ 2.1 Medical Visual Question Answering
The task of Med-VQA is to provide answers based on professional questions posed by the inquirer regarding medical images. In terms of training paradigms, early works[16][15][17] mostly employed RNNs and CNNs to respectively extract textual and visual features. However, these models often suffer from poor generalization. Thanks to the application of transformers, large-scale pretraining has begun to migrate from the textual domain to the multimodal domain. Pretraining VLP models[5] [6] using image-text pairs and fine-tuning them for downstream tasks has become the preferred approach for most multimodal tasks.
In terms of content output, previous works in Med-VQA have followed the VQA paradigm in the natural image domain, treating VQA as a classification task[16][15][17][18][19]. Specifically, fully connected layers and softmax layers are installed at the output end of the model to calculate cross-entropy loss for all candidate answers. Recently, some works[12] have also employed text-based decoders, which calculate masked language model (MLM) loss for all candidate answers and select the answer with the smallest loss as the model’s output. This approach is referred to as answer ranking. Although these methods achieve good accuracy on some benchmarks, they still treat visual question answering as a simple multi-classification task. When transferred to practical tasks without candidate answer pools, these Med-VQA methods cannot be effective.
In this paper, we propose a pretraining-finetuning paradigm called MISS for Med-VQA tasks. To our knowledge, this is the first work in the medical field that fully treats visual question answering as a text-generation task.
@@ 2.2 Visual-Language Pretraining Dataset
Currently, some works train Med-VQA models with the pretraining-finetuning paradigm. However, the medical field faces a shortage of image-text pairs for pretraining. ROCO[20] collects a large-scale unimodal and multimodal medical dataset from PubMed Central articles and constructs an image-text pairs dataset containing multiple types of medical images by expert radiologists. MediCaT[22] extracts images and corresponding captions from 131k openly available biomedical papers to construct a dataset containing more than 217k medical images with corresponding captions. However, these methods are similar to those used in the natural image domain to construct multimodal datasets by extracting news images and titles, and the collected images and captions contain a lot of noise, such as citations, labels, and other irrelevant information. Other high-quality open-source medical images for tasks such as classification and segmentation have been ignored because annotating these images also requires high costs and professional knowledge. We propose an automatic method for generating captions for unimodal images. This pioneering method attempts to utilize LLMs to construct a multimodal medical image-text dataset. The image data is clean, and the captions conform to human expression habits.
@@ 3.1 Overview
We adopt the pretraining and fine-tuning paradigm for training medical VQA models. In the pretraining stage, we first use image-text pairs to enable the model to learn multi-modal feature representation. In the fine-tuning stage, we use image-question pairs to train the model, enabling it to be applied to Med-VQA tasks ultimately. In the following, we will first introduce our model structure, followed by our pretraining method. Finally, we will present the TransCap method and the implementation details of fine-tuning.
3.2 Model Articture
@@
Our model adopts the encoder-decoder structure in its entirety. Unlike most dual-tower image-text VLP models in the past, our model’s encoder is divided into only two parts - the image encoder and the JTM encoder. For the image encoder, we borrow from the settings adopted by most recent works, which utilize the visual transformer as the image feature extractor. For an image input $I$, it’s firstly reshaped into flattened 2D patches and then encoded into a embedding sequence $\{x_{<cls>},x_{1},...,x_{n}\}$. After that, a 12-layer transformer encoder will extract its high-dimensional features.
@@
The JTM encoder is based on Bert and replaces the text and multi-modal encoders used in recent works. It performs representation learning of text and multi-modal features simultaneously. As shown in Figure 1(a), each JTM encoder is composed of 12 transformer-based layers, with each layer containing a bidirectional self-attention layer, a cross-attention layer, and a feed-forward layer. For each text input $T$, it’s first pre-processed by the tokenizer into a token sequence. Then, we feed it into the JTM encoder for multi-layer representation learning, where it interacts with the image features through cross-attention. Specifically, we define the text representation as $\{w_{<cls>},w_{1},...,w_{n}\}$, and the image embeddings are defined as $\{v_{<cls>},v_{1},...,v_{n}\}$. Both of these representations fuse and compute multimodal representations through $CrossAttention(Q,K,V)=SoftMax(\frac{QK^{T}}{\sqrt{d}}+B)V$, where text representation $\{w_{<cls>},w_{1},...,w_{n}\}$ generates query vectors $Q$, and the image representation $\{v_{<cls>},v_{1},...,v_{n}\}$ generates key $K$ and value vectors $V$.
@@
The decoder part of the model includes a text decoder, which aims to decode the multi-modal feature representation obtained by the JTM encoder into an output text representation. The backbone of the text decoder is similar to that of the JTM encoder but replaces the bidirectional self-attention layer in the JTM’s per-layer with a causal-attention layer. The text input passes through the causal-attention layer to calculate the text feature representation and then undergoes feature interaction with the multi-modal features through the cross-attention layer. The final features obtained are decoded by the tokenizer to obtain the text output.
3.3 Pretraining
@@
The VLP aims to align the multimodal features while trying to make the image encoder understand the feature distribution of images in high-dimensional space and comprehend the deep semantic information of medical images. Inspired by METER, we choose Image-Text Contrastive Learning (ITC), Image-Text Matching (ITM) and Mask Language Modeling (MLM) tasks for multi-modal pretraining.
@@
To enable the JTM encoder to learn the joint representation of text-multimodal features without being disturbed by the flow of features from another modality during the process of learning one representation, we adopt the method of BLIP[2] and deform the layer structure of the JTM encoder in different pretraining tasks. Specifically, at the beginning of model pretraining, the distance between $\{v_{<cls>},v_{1},...,v_{n}\}$ extracted by the image encoder and $\{w_{<cls>},w_{1},...,w_{n}\}$ of the JTM encoder in high-dimensional space is too far. At this time, it’s difficult for the two features to interact and perform ITM and MLM training. Therefore, at the beginning of training, the JTM encoder will discard the cross-attention layer and extract word embeddings so that narrowing the distance between $\{v_{<cls>},v_{1},...,v_{n}\}$ and $\{w_{<cls>},w_{1},...,w_{n}\}$ in high-dimensional space through the ITC task. The ITC, ITM, and LM losses are calculated as delineated below.
@@
Image-Text Contrastive Learning aims to learn unimodal representations before fusion[1]. ITC loss measures the distance of two embeddings in the feature space by a matrix similarity measure $\mathbf{S}=A^{T}B$. Inspired by MoCo, two momentum encoders are created and they respectively have the same architecture as the text encoder and the JTM encoder. Two queues are constructed to store the most recent $M$ image-text representations. The image and text features extracted by the image encoder and JTM encoder are denoted as $e_{V}(v_{cls})$ and $e_{J}(t_{cls})$, and those extracted by momentum encoders are denoted as $e^{\prime}_{V}(v^{\prime}_{cls})$ and $e^{\prime}_{J}(t^{\prime}_{cls})$. So we can calculate similarity $\mathbf{S}(I,T)=e_{V}(v_{cls})^{T}e^{\prime}_{J}(t^{\prime}_{cls})$ and $\mathbf{S}(T,I)=e_{J}(t_{cls})^{T}e^{\prime}_{V}(v^{\prime}_{cls})$, the softmax-normalized similarity between each image-text is calculated as follows:
$\small\displaystyle{p_{m}^{I2T}=\frac{exp(\mathbf{S}(I,T_{m})/\tau}{\Sigma_{m=1}^{M}exp(\mathbf{S}(I,T_{m})/\tau},p_{m}^{T2I}=\frac{exp(\mathbf{S}(T,I_{m})/\tau}{\Sigma_{m=1}^{M}exp(\mathbf{S}(T,I_{m})/\tau}}$ (1)
@@
$\small\displaystyle{\mathcal{L}=\frac{1}{2}\mathbb{E}_{(I,T)\sim D}[H(y^{I2T}(I),p^{I2T}(I))+H(y^{T2I}(T),p^{T2I}(T)]},$ (2)
@@
Mask Language Modeling Learning trains the text decoder by randomly masking some tokens in the word vectors. Unlike most VLP models that adopt text decoders that only receive multi-modal features, our decoder simultaneously accepts input from the original tokenized text and the JTM encoder. As shown in Figure 1(a), after the word vectors $\{w_{<decod>},w_{1},...,w_{n}\}$ undergo casual attention to extract word embeddings, they serve as query vectors $Q$ and interact with image embeddings which generate key $K$ and value vectors $V$ to calculate cross-attention. The MLM loss $\mathcal{L}_{MLM}$ is calculated similarly to that adopted in Bert, the details of which will be covered in the Appendix.
3.4 Tranfer and Caption

### v3

# MISS: A Generative Pre-training and Fine-tuning Approach for Med-VQA

@@ Abstract
Medical visual question answering (VQA) is a challenging multimodal task, where Vision-Language Pre-trained (VLP) models can effectively improve the generalization performance. However, most current methods in the medical field treat VQA as an answer classification task which is difficult to transfer to practical application scenarios. Additionally, due to the privacy of medical images and the expensive annotation process, large-scale medical image-text pairs datasets for pretraining are severely lacking. In this paper, we propose an efficient MultI-task Self-Supervised-learning-based framework (MISS) for medical VQA tasks. Unlike existing methods, we treat medical VQA as a generative task. We unify the text encoder and multimodal encoder and align image-text features through multi-task learning. Furthermore, we propose a Transfer-and-Caption (TransCap) method that extends the feature space of single-modal image datasets using Large Language Models (LLMs), enabling those traditional medical vision-field task data to be applied to VLP models. We conduct extensive experiments and compare them with existing medical VQA methods adopting a no-generative paradigm. We demonstrate the advantages of pre-training with data generated by the TransCap method and our method achieves excellent results with fewer multimodal datasets. The code has been released at https://github.com/TIMMY-CHAN/MISS.git.
Keywords:
Medical visual question answering Vision-language pre-training Multi-modal learning
@@ 1 Introduction
Visual Question Answering (VQA) is a multi-modal task based on vision and language, aiming to provide corresponding answers to the given images and questions. Thanks to the development of Convolutional Neural Networks (CNN) and Natural Language Processing (NLP) techniques, some works[23, 6, 10] have attempted to use CNN and Recurrent Neural Networks (RNN) to extract image and text features respectively for VQA tasks. With the emergence of transformers[28], image features and text features can be more easily embedded into the feature space with the same dimension, and VLP models[19, 26] have emerged continuously and have been proven to be effective solutions for downstream multi-modal tasks.
While effective, these VLP models have several limitations when applied to medical fields.
Compared with the VQA of natural images, Medical VQA (Med-VQA) requires a deeper and more accurate understanding of medical images. At the same time, due to the privacy of medical images and the high cost of high-quality text annotation, large-scale datasets for training Med-VQA are extremely scarce. Therefore, currently, Med-VQA is still a highly challenging task.
Currently, some multimodal models specialized in the medical domain have been proposed, such as M3AE [3], MRM [29], and CMITM [2], which unify Masked AutoEncoders (MAE) [11] and Masked Language Modeling (MLM) pre-training to learn joint representations of images and texts; MUMC [20] utilizes Masked Image Modeling(MIM) by sending masked images to the image encoder as data augmentation; PMC-CLIP builds a new large medical dataset and trains it on a CLIP [26] which pre-trained on natural images. However, the above Med-VQA models still have two key problems:
@@
b. They utilize image-text pairs crawled from the article centres for pre-training, which contain a lot of noise, and high-quality open-source medical images for other tasks, such as medical image classification and segmentation, have been being ignored.
In this paper, we propose a new pre-training and fine-tuning paradigm for medical image-text tasks, named MISS, and apply it to the Med-VQA task. We treat Med-VQA as an answer-generating task, making our method directly applied to real-world scenarios and generating responses that more closely match human expression. Unlike previous dual-tower multi-modal models, We innovatively unify the text encoder and multi-modal encoder, building a Joint Text-Multimodal (JTM) encoder and enabling it to learn joint feature representations using a multi-task learning approach.
@@
$\bullet$ We introduce a new pre-training and fine-tuning framework named MISS. Not considering the Large-scale Vision-Language models (parameters more than 1B), it is the first pure generative VQA model in the medical field.
@@ 2.1 Medical Visual Question Answering
The task of Med-VQA is to provide answers based on professional questions posed by the inquirer regarding medical images. In terms of training paradigms, early works[23, 6, 10, 14] mostly employed RNNs and CNNs to respectively extract textual and visual features. However, these methods often suffer from poor generalization. Thanks to the application of transformers, large-scale pretraining has begun to migrate from the textual domain to the multimodal domain. Training VLP models[19, 26] using image-text pairs and fine-tuning them for downstream tasks has become the preferred approach for most multimodal tasks.
In terms of content output, previous works in the Med-VQA field have followed the VQA paradigm in the natural image domain, treating VQA as a classification task[23, 6, 10, 21, 9]. Specifically, fully connected layers and softmax layers are installed at the output end of the model to calculate cross-entropy loss for all candidate answers. Recently, some works [20] have also employed text-based decoders, which calculate MLM loss for all candidate answers and select the answer with the smallest loss as the model’s output. This approach is referred to as answer ranking. Although these methods achieve good accuracy on some benchmarks, they still treat VQA as a simple multi-classification task. When transferred to practical tasks without candidate answer pools, these Med-VQA methods cannot be effective.
In this paper, we propose a pretraining-finetuning paradigm called MISS for Med-VQA tasks. To our knowledge, for small-scale VLMs, this is the first work in the medical field that fully treats VQA as a text-generation task.
@@ 2.2 Visual-Language Pretraining Dataset
Currently, many work train Med-VQA models with the pretraining-finetuning paradigm. However, the medical field faces a shortage of high-quality image-text pairs for pretraining. ROCO [25] collects a large-scale unimodal and multimodal medical dataset from PubMed Central articles and constructs an image-text pairs dataset containing multiple types of medical images by expert radiologists. MediCaT[27] extracts images and corresponding captions from 131k openly available biomedical papers to construct a dataset containing more than 217k medical images with corresponding captions. However, these methods are similar to those used in the natural image domain to construct multimodal datasets by extracting news images and titles from the internet, and the collected images and captions contain a lot of noise, such as citations, labels, and other irrelevant information. Other high-quality open-source medical images for tasks such as medical image classification and segmentation have been ignored because annotating these images also requires high costs and professional knowledge.
To deal with the above challenges, we propose an automatic method named TransCap for generating captions for unimodal images. This pioneering method attempts to utilize LLMs to construct a multimodal medical image-text dataset. The image data is clean, and the captions conform to human expression habits.
@@ 3 Method
Figure 1: Pretraining (a) and Finetuning (b) of our proposed method. We propose a pretraining and finetuning framework Miss for Med-VQA tasks which is composed of an image encoder, a JTM encoder, and a text decoder. ITC, ITM, and MLM Learning are used for pretraining. In the finetuning stage, the joint feature interacts with tokenized answers for MLM Learning.
@@ 3.1 Overview
We adopt the pre-training and fine-tuning paradigm for training medical VQA models. In the pretraining stage, we first use image-text pairs to enable the model to learn multi-modal feature representation. In the fine-tuning stage, we use image-question pairs to train the model, enabling it to be applied to Med-VQA tasks ultimately. In the following, we will first introduce our model structure, followed by our pretraining method. Finally, we will present the TransCap method and the implementation details of fine-tuning.
3.2 Model Architecture
@@
The architecture of MISS is demonstrated in Figure 1 a, our model adopts the encoder-decoder structure in its entirety. Unlike most dual-tower image-text VLP models in the past, our model’s encoder is divided into only two parts - the image encoder and the JTM encoder. For the image encoder, we borrow from the settings adopted by most recent works, which utilize the vision transformer(ViT)[7] as the image feature extractor. For an image input $I$, it’s firstly reshaped into flattened 2D patches and then encoded into a embedding sequence $\{x_{<cls>},x_{1},...,x_{n}\}$. After that, a 12-layer transformer encoder will extract its high-dimensional features.
@@
The JTM encoder replaces the text and multi-modal encoders used in recent works and it performs representation learning of text and multi-modal features simultaneously. As shown in Figure 1(a), each JTM encoder is composed of 12 transformer-based layers, with each layer containing a bidirectional self-attention layer, a cross-attention layer, and a feed-forward layer. For each text input $T$, it’s first pre-processed by the tokenizer into a token sequence. Then, we feed it into the JTM encoder for multi-layer representation learning, where it interacts with the image features through cross-attention. Specifically, we define the text representation as $\{w_{<cls>},w_{1},...,w_{n}\}$, and the image embeddings are defined as $\{v_{<cls>},v_{1},...,v_{n}\}$. Both of these representations fuse and compute multimodal representations through
$CrossAttention(Q,K,V)=SoftMax(\frac{QK^{T}}{\sqrt{d}}+B)V,$ (1)
where text representation $\{w_{<cls>},w_{1},...,w_{n}\}$ generates query vectors $Q$, and the image representation $\{v_{<cls>},v_{1},...,v_{n}\}$ generates key vectors $K$ and value vectors $V$.
@@
The decoder part of the model includes a text decoder, which aims to decode the multimodal feature representation obtained by the JTM encoder into an output text representation. The backbone of the text decoder is similar to that of the JTM encoder but replaces the bidirectional self-attention layer in the JTM’s per-layer with a causal-attention layer. The text input passes through the causal-attention layer to calculate the text feature representation and then undergoes feature interaction with the multi-modal features through the cross-attention layer. The final features obtained are decoded by the tokenizer to obtain the text output.
3.3 Pre-training
@@
The pre-training of Vision-Language Models (VLMs) aims to align the multimodal features while trying to make the image encoder understand the feature distribution of images in high-dimensional space and comprehend the deep semantic information of medical images. Inspired by METER[8], we choose Image-Text Contrastive Learning (ITC), Image-Text Matching (ITM) and Mask Language Modeling (MLM) tasks for multi-modal pretraining.
@@
To enable the JTM encoder to learn the joint representation of text-multimodal features without being disturbed by the flow of features from another modality during the process of learning one representation, we adopt the method of BLIP[17] and deform the layer structure of the JTM encoder in different pretraining tasks. Specifically, at the beginning of model pretraining, the distance between $\{v_{<cls>},v_{1},...,v_{n}\}$ extracted by the image encoder and $\{w_{<cls>},w_{1},...,w_{n}\}$ of the JTM encoder in high-dimensional space is too far. At this time, it’s difficult for the two features to interact and perform ITM and MLM training. Therefore, at the beginning of training, the JTM encoder will discard the cross-attention layer and extract word embeddings so that narrowing the distance between $\{v_{<cls>},v_{1},...,v_{n}\}$ and $\{w_{<cls>},w_{1},...,w_{n}\}$ in high-dimensional space through the ITC task. The ITC, ITM, and MLM losses are calculated as delineated below.
@@
Image-Text Contrastive Learning aims to learn unimodal representations before fusion[18]. ITC loss measures the distance of two embeddings in the feature space by a matrix similarity measure $\mathbf{S}=A^{T}B$. Inspired by MoCo[12], two momentum encoders are created and they respectively have the same architecture as the text encoder and the JTM encoder. Two queues are constructed to store the most recent $M$ image-text representations. The image and text features extracted by the image encoder and JTM encoder are denoted as $e_{V}(v_{cls})$ and $e_{J}(t_{cls})$, and those extracted by momentum encoders are denoted as $e^{\prime}_{V}(v^{\prime}_{cls})$ and $e^{\prime}_{J}(t^{\prime}_{cls})$. So we can calculate similarity $\mathbf{S}(I,T)=e_{V}(v_{cls})^{T}e^{\prime}_{J}(t^{\prime}_{cls})$ and $\mathbf{S}(T,I)=e_{J}(t_{cls})^{T}e^{\prime}_{V}(v^{\prime}_{cls})$, the softmax-normalized similarity between each image-text is calculated as follows:
${p_{m}^{I2T}=\frac{exp(\mathbf{S}(I,T_{m})/\tau}{\Sigma_{m=1}^{M}exp(\mathbf{S}(I,T_{m})/\tau},p_{m}^{T2I}=\frac{exp(\mathbf{S}(T,I_{m})/\tau}{\Sigma_{m=1}^{M}exp(\mathbf{S}(T,I_{m})/\tau}},$ (2)
@@
${\mathcal{L}=\frac{1}{2}\mathbb{E}_{(I,T)\sim D}[H(y^{I2T}(I),p^{I2T}(I))+H(y^{T2I}(T),p^{T2I}(T)]},$ (3)
@@
Mask Language Modeling Learning trains the text decoder by randomly masking some tokens in the word vectors. Unlike most VLP models that adopt text decoders that only receive multi-modal features, our decoder simultaneously accepts input from the original tokenized text and the JTM encoder. As shown in Figure 1(a), after the word vectors $\{w_{<decod>},w_{1},...,w_{n}\}$ undergo causal attention to extract word embeddings, they serve as query vectors $Q$ and interact with image embeddings which generate key $K$ and value vectors $V$ to calculate cross-attention. The MLM loss $\mathcal{L}_{MLM}$ is calculated similarly to that adopted in BERT[5], where 15% of the tokens are randomly selected. Then 80% of selected tokens are replaced with a special token [MASK], 10% are randomly replaced with other words, and the other 10% are left unchanged. In our proposed paradigm, while pretraining the model, the tokenized input embeddings $\{w_{<decod>},w_{1},...,w_{n}\}$ and the image embeddings $\{i_{<cls>},i_{1},...,i_{n}\}$ fuse in the cross-attention layer. Then finetune the joint feature $f$ who fuses with question embeddings $\{q_{<encod>},q_{1},...,q_{n}\}$. Both of them are then sent to the feed-forward layer and calculate the minimized cross-entropy loss between predicted results and ground-truth results:
${\mathcal{L}_{MLM}=\mathbb{E}_{(I,\hat{T})\sim D}H(y^{msk},p^{msk}(I,\hat{T})},$ (5)
3.4 Transfer and Caption